A firm has the following capital structure: £100 million of equity (market value) with 100 million shares outstanding, and £100 million of debt. The beta of the firm’s stock is 1.6. The firm’s cost of equity is 10 percent, and the yield on riskless bonds is 2 percent. There is no tax. Assuming that the firm can borrow at the risk-free rate and that both CAPM (Capital Asset Pricing Model) and the Modigliani-Miller theorem hold, answer the following questions. i) What is the share price of the firm? ii) What is the WACC of the firm? iii) What should be the expected return on the market portfolio? iv) Suppose the firm changes its capital structure so that its debt increases to £140 million, and the equity decreases by £60 million. What should be the firm’s cost of equity after the change?
Answer #1

(i) Equity Value =  £100 million and Number of Shares Outstanding = 100 million

Share Price = 100 / 100 =  £1 per share

(ii) Cost of Equity = 10 %, Cost of Debt = Firm's Borrowing Rate = Risk-Free Rate = Yield on Riskless Bond = 2 %

Tax Rate = 0

WACC = 10 x (100 / 200) + 2 x (100/200) = 6 %

(iii) Let the market portfolio's return be Rm, Firm Beta = 1.6

As CAPM Holds, we have Cost of Equity = Risk-Free Rate + Firm Beta x (Rm - Risk-Free Rate)

10 = 2 + 1.6 x (Rm - 2)

Rm = 7 %

(iv) Let the unlevered cost of equity be Ru

Therefore, as per MM theory 10 = Ru + Debt-to-Equity Ratio x (Ru - Rd) where Rd is the cost of debt

10 = Ru + 1 x (Ru - 2)

Ru = 6 %

As new Capital Structure has debt worth  £140 million and equity worth  £60 million, the new Debt-to-Equity Ratio becomes (140 / 60) = 7/3

Therefore, new cost of equity = Ru + New Debt-to-Equity Ratio x (Ru - Rd)

Cost of Equity = 6 + (7/3) x (6 - 2) = 15.33 %